OpenStudy (anonymous):

Check my math please :) Rationalize the denominator: sqrt{a}-sqrt{b}/sqrt{a}+sqrt{b}

OpenStudy (anonymous):

gr! \[\sqrt{a}-\sqrt{b}/\sqrt{a}+\sqrt{b}\]

OpenStudy (anonymous):

I got: \[a-2\sqrt{ab}-b/a+b\]

OpenStudy (saifoo.khan):

\[\frac{\sqrt{a}-\sqrt{b}}{\sqrt{a}+\sqrt{b}}\]

OpenStudy (anonymous):

Hey Saifoo, how do you get the fractions? Cant see them in the equation editor.

jimthompson5910 (jim_thompson5910):

should be \[\LARGE \frac{a-2\sqrt{ab}-b}{a-b}\]

OpenStudy (saifoo.khan):

\frac{kd}{\ s im}

OpenStudy (anonymous):

Thanks Jim, almost had it :)!

OpenStudy (saifoo.khan):

like that.

OpenStudy (saifoo.khan):

Jim's right.

OpenStudy (anonymous):

Thanks Saifoo too!

OpenStudy (saifoo.khan):

Np.

jimthompson5910 (jim_thompson5910):

if you right click any image, you'll get the source for instance if you right click the last image I posted, you'll get \frac{a-2\sqrt{ab}-b}{a-b}

jimthompson5910 (jim_thompson5910):

that will tell you exactly how it was entered

OpenStudy (anonymous):

awesome!